# Rasterizing Wireless Radiance Field via Deformable 2D Gaussian Splatting

### v2

@@ 1 Introduction
In modern communication systems, digital information is conveyed through radio frequency (RF) waveforms that interact with the physical environment. As they propagate, these signals experience reflection, diffraction, scattering, and absorption, resulting in a continuous and spatially varying signal field (see Fig. 1). Characterizing the spatial–frequency structure of this field forms the foundation of wireless radiance field (WRF) modeling [1], a key area in advancing wireless communication. The motivation for recovering the WRF lies in its ability to encapsulate rich geometric and spectral information, which in turn supports channel estimation, as well as downstream tasks including localization and environmental sensing. Over the years, researchers have explored a wide range of techniques to tackle this problem, from classical Maxwell-based formulations to modern deep neural network approaches. As 6G technologies continue to advance, communication systems in latency-critical domains such as the Internet of Things (IoT) are expected to deliver ultra-reliable, low-latency performance, with end-to-end delays ranging from 10 to 100 microseconds [2, 3]. This makes fast and accurate WRF modeling essential for the practical deployment of next-generation wireless systems.
Figure 1: A toy example of wireless radiance field.
@@
Motivation: Existing approaches to WRF modeling can be broadly categorized into two types: rule-based and learning-based [4]. Rule-based techniques are typically divided into probabilistic and deterministic models. The former rely on empirical formulas calibrated from measurement data to estimate signal strength as a function of transceiver location. While this allows for fast computation, such models often fail to capture fine-grained spatial variations and lack reconstruction fidelity [5, 6]. In contrast, deterministic techniques like ray tracing simulate RF propagation using physical laws, taking into account environmental geometry and material properties. These methods can achieve high accuracy, but their effectiveness hinges on the availability of detailed 3D scene models and comes at a significant computational cost [7]. To overcome the limitations of rule-based methods, learning-based approaches have been developed to reconstruct WRF spectra directly from data using deep neural networks. Classical models such as variational autoencoders (VAEs) [8] and generative adversarial networks (GANs) [9] aim to learn a low-dimensional continuous manifold of the WRF from observed spectra. Once trained, these models can interpolate spectra at arbitrary positions to recover the full field. Despite their improved reconstruction quality, the training process often requires large amounts of high-quality data and offers limited interpretability due to their black-box architectures.
Emerging learning-based studies have drawn connections between WRFs and optical radiance fields, recognizing that both can be modeled as continuous spatial functions based on radiative propagation. This conceptual equivalence has motivated the adaptation of novel view synthesis (NVS) techniques to WRF modeling, where synthesizing spectra at unseen positions is analogous to rendering novel views [10, 11]. A representative example in the NVS domain is neural radiance fields (NeRF) [12], which reconstructs the optical radiance field from sparse views via multilayer perceptron (MLP)-based volumetric rendering. Building on this idea, [13, 14, 15] extend the NeRF framework to the wireless domain, achieving impressive spectra reconstruction fidelity by implicitly modeling the underlying WRF structure. However, their accuracy comes at the cost of expensive MLP queries for field density, which significantly slows down inference and limits their suitability for latency-sensitive applications.
Recent advances in NVS have seen Gaussian splatting (GS) [16] achieves notable progress in both efficiency and rendering quality. Unlike NeRF-based methods that rely on neural volumetric rendering, GS models a scene as a set of anisotropic 3D Gaussian primitives with explicit geometric and radiance attributes. These primitives collectively approximate the underlying optical radiance field and render novel views via differentiable rasterization. A key strength of the GS pipeline lies in its CUDA-accelerated parallel implementation, which enables differentiable rasterization to be efficiently performed in a tile-wise manner. By avoiding computationally expensive volumetric ray marching, GS significantly reduces both training and inference overhead while preserving photorealistic rendering. Motivated by its efficiency in modeling optical radiance fields, we aim to adopt the GS technique to tackle the remaining challenges in WRF. Yet, our analysis reveals several limitations in its direct application to the WRF domain, as outlined below to motivate our redesign.
@@
Observation: To our knowledge, existing GS-based methods [4, 17] have been applied primarily to WRF scenarios with one moving transmitter (TX) and one fixed receiver (RX). At each TX position, the RX records a spatial spectrum synthesized via differentiable rasterization of a set of 3D Gaussians, with their attributes dynamically predicted from the TX’s location. While this design improves inference efficiency, we identify several limitations that restrict its applicability in the wireless domain: First, the use of a 3D volumetric representation is unnecessary. The 3D Gaussian formulation captures spatial variations critical in NVS, but such variations are absent in WRF under TX-side mobility. In this setting, TX movement induces only a global shift of the 3D Gaussian splatting (3DGS) field, without introducing any spatial variation in its projection from the RX side. As the projection remains constant, we posit that modeling the 2D projection of 3D Gaussian directly may serve as an equally effective solution. Second, existing methods lack locality-aware modeling. In [4], Gaussian attributes are predicted independently at each TX location, without accounting for spatial correlations among neighboring positions [4]. This position-wise formulation overlooks the continuity of the underlying physical field, potentially leading to suboptimal convergence and poor generalization [18]. Third, sparse supervision introduces a risk of overfitting. Existing methods optimize quality loss only over observed regions, without applying any regularization to unobserved areas. This lack of supervision hinders the model’s ability to interpolate and generalize to unseen positions. Fourth, limited evaluation settings. Existing methods are typically tested on a single scene under the TX-moving configuration ($\text{NeRF}^{2}-\text{s}23$ dataset [13]). The lack of diverse scenes, i.e., RX-moving cases, limits the credibility of their results.
Together, these limitations constrain the scalability, generalization, and physical alignment of existing GS-based WRF modeling approaches. In response, we propose a streamlined GS framework tailored to the structural and physical characteristics of the WRF, and enabling evaluation across both TX- and RX-moving scenarios.
@@
Our method: We propose a deformable 2D Gaussian splatting (2DGS) framework, termed SwiftWRF, to address WRF reconstruction under single-sided transceiver [^1] Refers to either the moving TX or RX, depending on the scenario. mobility. The method begins by formulating a 2D Gaussian representation tailored to the wireless domain, where each primitive captures localized signal behavior through a 2D Gaussian distribution and a complex-valued response. To render the spatial spectrum, we adopt a Gaussian mixture model (GMM)–like formulation [19], in which all 2D Gaussian primitives are directly blended over the plane via differentiable rasterization. Compared to 3D Gaussians, our 2D formulation substantially improves compactness, reducing the parameter count by a factor of seven, as illustrated in Fig. 2. The blending process also eliminates the need for camera projection and removes dependence on geometry-based initialization. Complementing this, transceiver mobility is modeled through a continuous deformation field, implemented as a lightweight MLP that predicts position-dependent offsets to a global set of 2D Gaussian parameters. This design enables fast, spatially coherent adaptation across locations, aligning more closely with the physical characteristics of WRF. To improve generalization and mitigate overfitting, we incorporate an annealed smoothing mechanism that injects controlled positional noise during training, with its magnitude decaying exponentially as training progresses.
For evaluation, we construct a synthetic dataset to assess the generality of SwiftWRF beyond TX-moving settings. The dataset includes both TX-moving and RX-moving scenarios, spanning three representative indoor layouts, with spectra generated using MATLAB’s ray-tracing toolbox. Extensive experiments show that SwiftWRF achieves a peak signal-to-noise ratio (PSNR) improvement of 2.33–10.41 dB over existing baselines, while delivering inference speeds of approximately 100k FPS—over 500$\times$ faster than WRF-GS and more than 20,000$\times$ faster than NeRF2. Finally, we evaluate SwiftWRF on two case studies. For angle-of-arrival (AoA) prediction, the synthesized spectra are integrated into a turbo-learning framework, where they serve as augmented data to enhance estimation accuracy. For received signal strength indicator (RSSI) prediction, we adapt the rendering head with average pooling to support scalar field regression. These case studies validate the generalizability of our method and demonstrate consistent improvements over existing baselines.
@@
  • We construct and release a synthetic WRF dataset covering both TX-moving and RX-moving scenarios, generated using MATLAB’s ray-tracing toolbox over multiple representative indoor environments.
@@ 2.1 Wireless Radiance Field Modeling
WRF modeling seeks to characterize electromagnetic wave propagation between TXs and RXs, incorporating reflection, diffraction, scattering, and path‐loss effects [20, 21]. Conventional approaches to channel modeling can be broadly classified into two categories: probabilistic models and deterministic ray-tracing models. The former are computationally efficient and easy to implement; however, they often fail to capture fine-grained spatial variations and site-specific details [22]. The latter offer high-fidelity, site-specific predictions by modeling the exact propagation environment [1], but they are computationally intensive and require detailed prior knowledge of the geometric layout, which limits their scalability and applicability in dynamic or unknown environments. Learning‐based methods instead train deep networks on extensive, location‐tagged channel measurements to directly infer complex propagation phenomena [8, 9], but they often lack interpretability and struggle to generalize beyond their training distributions. More recently, NeRF–based methods have been applied to learn WRF as an optical field [13, 14, 23]. Zhao et al. [13] developed $\text{NeRF}^{2}$ for spatial‐spectrum reconstruction in real environments; Orekondy et al. [23] introduced a NeRF surrogate for indoor EM propagation using synthetic datasets; and Lu et al. [14] proposed NeWRF, which predicts channel states from sparse measurement sets. Considering the slow rendering of NeRF, a few GS–based methods model WRF with explicit Gaussian primitives and leverage CUDA-accelerated rasterization to synthesize spectra in near real time [17, 4]. However, these approaches remain rudimentary 3D GS frameworks and lack the specialized design needed to meet the stringent requirements of ultra-reliable low-latency communication (URLLC). To address this gap, we present SwiftWRF, an efficient 2DGS framework that simultaneously achieves faster training and inference times, alongside superior reconstruction fidelity.
@@ 3.1 Wireless Radiance Field
[…]
Assuming $L_{p}$ multipath components, the channel at $R_{m,n}$ is
$h_{m,n}=\sum_{l=1}^{L_{p}}\rho_{l}e^{j\Delta\alpha_{m,n}(\theta_{l},\phi_{l})},$ (2)
where $\rho_{l}$ is the power coefficient associated with the $l$-th propagation path, and $\theta_{l}$, $\phi_{l}$ denote its azimuth and elevation angles, respectively. The phase difference is expressed as
$\Delta\alpha_{m,n}(\theta,\phi)$ $=\alpha_{m,n}(\theta,\phi)-\alpha_{1,1}(\theta,\phi)$
$=2\pi\Delta d_{m,n}(\theta,\phi)/\lambda\mod 2\pi$
$=-2\pi r_{m,n}\cos(\theta-\beta_{m,n})\cos(\phi)/\lambda\mod 2\pi.$ (3)
Here, $\Delta d_{m,n}$ is the path length difference between antennas $R_{m,n}$ and $R_{1,1}$ along direction $(\theta,\phi)$, and $\lambda$ denotes the signal wavelength. The RX estimates the spatial spectrum by applying an analog combining matrix that steers beams toward predefined azimuth $\theta$ and elevation $\phi$ angles. When a beam aligns with the true AoA, signals from that direction combine constructively, while signals from other directions interfere destructively and are attenuated [25]. The spatial spectrum at direction $(\theta,\phi)$ is then computed as
@@
where $w_{m,n}(\theta,\phi)=e^{-j\Delta\alpha_{m,n}(\theta,\phi)}$ is the complex weight for steering a beam to a certain angle of $(\theta,\phi)$ applied to the $(m,n)$-th antenna element. Under our assumption of $1^{\circ}$ angular resolution, the spatial spectrum is represented as a $360\times 90$ matrix $\mathbf{A}$, defined by
@@
3.2 Problem Formulation
We consider the problem of WRF reconstruction under single-sided transceiver mobility, where exactly one TX and one RX are placed in an indoor scene. In this setting, either the TX or the RX moves while the other remains fixed, giving rise to two distinct scenarios:
TX-moving scenario. The TX moves across a set of known positions
$\mathcal{P}=\{\boldsymbol{p}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N_{s}},$
RX-moving scenario. The TX emits WRF signals from a fixed position $\boldsymbol{p}\in\mathbb{R}^{3}$, while the RX moves across a set of known positions
$\mathcal{Q}=\{\boldsymbol{q}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N_{s}},$
$\{(\boldsymbol{s}_{i},\mathbf{A}_{i})\}_{i=1}^{N_{s}},\quad\boldsymbol{s}_{i}\in\begin{cases}\mathcal{P},&\text{TX-moving},\\ \mathcal{Q},&\text{RX-moving},\end{cases}$
Figure 4: Overview of SwiftWRF. For illustration, we present the operating pipeline for the TX-moving scenario (top). The bottom left depicts the architecture of deformable 2DGS, while the bottom right illustrates the spectrum rasterization process.
@@ 3.3 Gaussian Splatting
3D Gaussian: 3DGS reconstructs unseen views of a 3D scene by training a set of Gaussian primitives from a sparse collection of captured images. These primitives are positioned in free 3D space, with each representing a small, smooth region characterized by a set of learnable attributes: a mean position $\boldsymbol{\mu}\in\mathbb{R}^{3}$, a covariance matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$, an opacity value $\alpha\in[0,1]$ indicating transparency, and a color vector $\boldsymbol{c}\in\mathbb{R}^{k}$ encoding appearance via spherical harmonics (SH) coefficients. Each Gaussian primitive models the likelihood of a point $\boldsymbol{x}\in\mathbb{R}^{3}$ belonging to its region, following a 3D Gaussian distribution defined as
@@
where $\boldsymbol{d}_{n}=\boldsymbol{x}_{i}-\boldsymbol{\mu}_{n}$ denotes the displacement between the pixel location $\boldsymbol{x}_{i}$ and the center of the $n$-th Gaussian.
3.4 Summary
To summarize, we describe the formulation of spatial spectrum and define our problem setup under single-sided transceiver mobility. We also outline the GS technique, particularly its 2D extension, which forms the basis of our proposed framework. The observation that WRF spectra share the same 2D structural nature as images motivates our choice of 2D Gaussians. Building on this formulation, the next section addresses the modality gap between images and spectra, and details how our framework incorporates transceiver mobility.

### v3

# Deformable 2D Gaussian Splatting for Efficient Wireless Radiance Field Rendering

@@ 1 Introduction
In modern communication systems, digital information is conveyed through radio frequency (RF) waveforms that interact with the physical environment. As these signals propagate from a transmitter (TX) to a receiver (RX), they undergo reflection, diffraction, scattering, and absorption, giving rise to a continuous electromagnetic signal field (see Fig. 2). To understand and utilize this complex propagation behavior, Wireless Radiance Field (WRF) modeling has emerged as a means to learn how the signal field’s frequency response varies across space [2]. The task is to reconstruct the spatial spectrum, which captures the omnidirectional signal distribution observed at an RX, along novel TX–RX deployments, using spectra collected from known ones. WRF modeling encapsulates rich geometric and spectral information that benefits downstream tasks such as channel estimation, localization, and environmental sensing. Over the years, researchers have explored a wide range of techniques to tackle this problem, from classical Maxwell-based formulations to modern deep neural network approaches. As 6G technologies continue to advance, communication systems in latency-critical domains such as the Internet of Things (IoT) are expected to deliver ultra-reliable, low-latency performance, with end-to-end delays ranging from 10 to 100 microseconds [24, 1, 43]. This makes fast and accurate WRF modeling essential for the practical deployment of next-generation wireless systems.
@@
Background: WRF modeling aims to represent how spatial spectra vary with transceiver location [19, 42]. Traditional rule-based approaches, including empirical path-loss models and ray tracing, offer a speed-accuracy trade-off: empirical models are fast but often miss fine-grained spatial variations, while ray tracing can be accurate but depends on detailed scene models and is computationally expensive [35, 11, 29, 30, 7]. Learning-based methods improve fidelity by leveraging deep generative models (e.g., variational autoencoders (VAEs) and generative adversarial networks (GANs)) to synthesize unobserved spectra from measurements [13, 28]. A recent line of work draws a direct analogy between WRF reconstruction and novel view synthesis (NVS), treating spectrum prediction at unseen locations as rendering from a neural radiance field (NeRF) [40, 46]. NeRF-style wireless methods adopt multilayer perceptron (MLP)-based volumetric rendering and achieve strong quality [23, 46, 20, 5], yet their inference remains expensive because each query requires many MLP evaluations along rays (or dense samples) and the cost grows with the target resolution and the number of queried locations. This limits practicality in latency-sensitive settings.
Figure 2: A toy example illustrating the wireless radiance field from a given TX-RX deployment.
Gaussian splatting (GS) offers a more efficient rendering paradigm by representing the field with explicit Gaussian primitives and computing outputs via CUDA-parallel rasterization [12, 6]. Compared with volumetric rendering, GS avoids repeated network queries at inference and instead aggregates contributions from a set of differentiable primitives in a tile-wise manner, which is well suited to real-time synthesis. These properties motivate us to adopt a GS-style explicit representation for WRF, so that spectra at novel positions can be reconstructed with high throughput while preserving reconstruction quality.
@@
Observation: While GS represents the state-of-the-art (SOTA) in field reconstruction, directly applying it to WRF modeling exposes several mismatches between the assumptions of the vision domain and the wireless domain. These gaps hinder the efficiency and effectiveness of GS in this context. First, 3D volumetric representation is redundant in the WRF setting. The WRF modeling task is typically defined as synthesizing spatial spectra at a fixed RX, where a moving TX emits distinct WRFs from different positions. In this setting, each WRF is modeled using a separate set of 3D Gaussians, followed by a constant mapping to form the RX’s spectrum. Since the RX’s position and orientation remain unchanged, only the components of the 3D Gaussians aligned with its view contribute to the observed signal (see Fig. 1 in the middle). This causes the 3D representation to degenerate into 2D Gaussians, rendering the volumetric design unnecessary. Second, existing methods lack locality-aware modeling. In [35], 3D Gaussian attributes are predicted independently at each TX location. However, as the TX follows a continuous trajectory and wireless propagation is governed by deterministic multipath effects, the resulting WRF varies smoothly in space. This implies strong correlations between spectra observed at neighboring TX locations. Third, sparse supervision introduces a risk of overfitting. Existing methods optimize quality loss only over observed regions, without applying any regularization to unobserved areas. This lack of supervision hinders the model’s ability to interpolate and generalize to unseen positions. Fourth, limited evaluation settings. Existing methods are typically tested on a single scene under the TX-side moving configuration ($\text{NeRF}^{2}-\text{s}23$ dataset [46]). The lack of diverse scenes, i.e., RX-side moving cases, limits the credibility of their results. Together, these limitations constrain the scalability, generalization, and physical alignment of existing GS-based WRF modeling approaches. In response, we propose a streamlined GS framework tailored to the structural characteristics of the WRF, and enabling evaluation across both TX-side and RX-side moving scenarios.
@@
Our method: We formulate the WRF modeling problem under a single-sided transceiver mobility setting, where the perceived WRF spectra between the TX–RX pair vary due to the motion of one moving side. This setting follows the evaluation protocol of NeRF2 and related WRF surrogates, but we depart from their MLP-based implicit field representation: instead of querying an MLP densely for rendering, we extend the explicit rasterization line of Gaussian splatting to the wireless spectrum domain. In particular, our representation inherits the 2D Gaussian rasterization idea used in image/video 2DGS works [45, 18], and adapts it to model antenna-array spectrum outcomes under mobility. To address this, we propose a deformable 2D Gaussian splatting (2DGS) framework, termed SwiftWRF. The method begins by formulating a 2D Gaussian representation tailored to the wireless domain, where each primitive captures localized signal behavior through a 2D Gaussian distribution and a complex-valued response. Unlike prior image/video 2DGS methods defined on the image plane, our primitives are formulated in the polar spectrum domain for antenna-array rasterization. We adopt a Cholesky-parameterized covariance and replace RGB appearance modeling with signal attenuation and a complex-valued response. To render the spatial spectrum, we adopt a Gaussian mixture model (GMM)–like formulation [45], in which all 2D Gaussian primitives are directly blended over the plane via differentiable rasterization. Compared to 3D Gaussians, our 2D formulation substantially improves compactness, reducing the parameter count by a factor of seven, as illustrated in Fig. 1. The blending process also eliminates the need for camera projection and removes dependence on geometry-based initialization. Complementing this, transceiver mobility is modeled through a continuous deformation field, implemented as a lightweight MLP that predicts position-dependent offsets to a global set of 2D Gaussian parameters. This design enables fast and spatially coherent adaptation across locations, effectively translating the spatial deformation of the WRF into a planar deformation in the 2D Gaussian space. For model training, we adopt a coarse-to-fine training protocol: the model is first trained without deformation on a shared set of 2D Gaussians, and subsequently fine-tuned with the deformation field enabled. To further improve generalization and mitigate overfitting, we introduce an annealed smoothing mechanism that injects controlled positional noise during fine-tuning. The noise magnitude decays exponentially over time, promoting smoother optimization in the early stages and higher reconstruction quality in the later stages.
[…]
@@
  • We construct and release a synthetic WRF dataset covering both TX-side moving and RX-side moving scenarios, generated using MATLAB’s ray-tracing toolbox over multiple representative indoor environments.
@@ 2.1 Wireless Radiance Field Modeling
WRF modeling seeks to characterize electromagnetic wave propagation between TXs and RXs, incorporating reflection, diffraction, scattering, and path‐loss effects [26, 34]. Conventional approaches to channel modeling can be broadly classified into two categories: probabilistic models and deterministic ray-tracing models. The former are computationally efficient and easy to implement; however, they often fail to capture fine-grained spatial variations and site-specific details [41, 15]. The latter offer high-fidelity, site-specific predictions by modeling the exact propagation environment [2], but they are computationally intensive and require detailed prior knowledge of the geometric layout, which limits their scalability and applicability in dynamic or unknown environments. Learning‐based methods instead train deep networks on extensive, location‐tagged channel measurements to directly infer complex propagation phenomena [13, 28], but they often lack interpretability and struggle to generalize beyond their training distributions. More recently, NeRF–based methods have been applied to learn WRF as an optical field [46, 20, 27, 3]. Zhao et al. [46] developed $\text{NeRF}^{2}$ for spatial-spectrum reconstruction in real environments; Orekondy et al. [27] introduced a NeRF surrogate for indoor EM propagation using synthetic datasets; Lu et al. [20] proposed NeWRF, which predicts channel states from sparse measurement sets; and EchoNeRF [3] learns a NeRF-style RF field from sparse RF measurements to infer indoor structure by modeling the received signal power as a combination of LoS and multipath reflection components. Considering the slow rendering of NeRF, several GS-based methods model WRF with explicit Gaussian primitives and leverage CUDA-accelerated rasterization for near real-time synthesis. RF-3DGS [44] reconstructs a radio radiance field and targets multimodal spatial-channel state information (CSI), e.g., gain/delay/AoA, by encoding channel attributes with CSI-aware spherical-harmonics representations, enabling fast CSI queries at arbitrary locations. RFSPM [37] focuses on scalable RF signal spatial propagation modeling, representing a scene with customized GPU kernels to efficiently predict received signal characteristics across diverse wireless settings. More recently, GSpaRC [25] pushes GS-based RF reconstruction toward real-time deployment by achieving sub-millisecond latency for RF channel/CSI reconstruction while maintaining competitive accuracy. However, these approaches remain rudimentary 3D GS frameworks and lack the specialized design needed to meet the stringent requirements of ultra-reliable low-latency communication (URLLC) [19, 17]. To address this gap, we present SwiftWRF, an efficient 2DGS framework that simultaneously achieves faster training and inference times, alongside superior reconstruction fidelity.
@@ 3.1 Wireless Radiance Field
$\left\{\begin{aligned} r_{m,n}&=D\sqrt{(m-1)^{2}+(n-1)^{2}},\\ \beta_{m,n}&=\arctan 2(m-1,n-1).\end{aligned}\right$. (1)
Assuming $L_{p}$ multipath components, the narrowband channel at $R_{m,n}$ is
$h_{m,n}=\sum_{l=1}^{L_{p}}\rho_{l}e^{j\Delta\alpha_{m,n}(\hat{\theta}_{l},\hat{\phi}_{l})},$ (2)
where $\rho_{l}$, $\hat{\theta}_{l}$, and $\hat{\phi}_{l}$ are the power coefficient, azimuth angle, and elevation angle, respectively, all associated with the $l$-th propagation path. According to [46], the phase difference is expressed as
$\Delta\alpha_{m,n}(\hat{\theta}_{l},\hat{\phi}_{l})$ $=\alpha_{m,n}(\hat{\theta}_{l},\hat{\phi}_{l})-\alpha_{1,1}(\hat{\theta}_{l},\hat{\phi}_{l})$
$=2\pi\Delta d_{m,n}(\hat{\theta}_{l},\hat{\phi}_{l})/\lambda\mod 2\pi,$ (3)
where $\Delta d_{m,n}(\hat{\theta}_{l},\hat{\phi}_{l})=-r_{m,n}\cos(\hat{\theta}_{l}-\beta_{m,n})\cos(\hat{\phi}_{l})$ represents the path length difference between antennas $R_{m,n}$ and the reference antenna $R_{1,1}$ along direction $(\hat{\theta}_{l},\hat{\phi}_{l})$, and $\lambda$ denotes the signal wavelength. The RX estimates the spatial spectrum by applying an analog combining matrix that steers beams toward predefined azimuth $\theta$ and elevation $\phi$ angles. When a beam aligns with the true AoA, signals from that direction combine constructively, while signals from other directions interfere destructively and are attenuated [4]. According to [46, 35], the spatial spectrum at direction $(\theta,\phi)$ is then computed as
@@
where $w_{m,n}(\theta,\phi)=e^{-j\Delta\alpha_{m,n}(\theta,\phi)}$ is the complex weight for steering a beam towards the angle $(\theta,\phi)$, applied to the $(m,n)$-th antenna element, and $\angle h_{m,n}$ denotes the phase of $h_{m,n}$. Under our assumption of $1^{\circ}$ angular resolution, the spatial spectrum is represented as a $360\times 90$ matrix $\mathbf{A}$, defined by
@@ 3.2 Gaussian Splatting
3D Gaussian: 3DGS reconstructs unseen views of a 3D scene by training a set of Gaussian primitives from a sparse collection of captured images [12]. These primitives are positioned in free 3D space, with each representing a small, smooth region characterized by a set of learnable attributes: a mean position $\boldsymbol{\mu}\in\mathbb{R}^{3}$, a covariance matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{3\times 3}$, an opacity value $\alpha\in[0,1]$ indicating transparency, and a color vector $\boldsymbol{c}\in\mathbb{R}^{k}$ encoding appearance via spherical harmonics (SH) coefficients. Each Gaussian primitive models the likelihood of a point $\boldsymbol{x}\in\mathbb{R}^{3}$ belonging to its region, following a 3D Gaussian distribution defined as
@@
where $\boldsymbol{d}_{n}=\boldsymbol{x}_{i}-\boldsymbol{\mu}_{n}$ denotes the displacement between the pixel location $\boldsymbol{x}_{i}$ and the center of the $n$-th Gaussian. 2D Gaussians were originally introduced for fast image/content reconstruction. In contrast, our goal is to predict wireless spectra at unseen positions, which requires both strategies to mitigate overfitting and a wireless-domain adaptation of the 2D Gaussian formulation to better capture signal-specific characteristics.
3.3 Problem Formulation
3.3.1 Problem Statement
We study the problem of reconstructing spatial spectra at unobserved transceiver positions using spectra collected from known locations, under a single-sided transceiver mobility setting. In this setting, either the TX or the RX moves in a multipath-rich environment, while both maintain fixed orientations. In the TX-side moving case, the WRF is dynamically emitted from the moving TX, while the mapping from the WRF to the observed spectrum remains constant due to a fixed RX. Given the constant field-to-spectrum mapping, the WRF variations are directly translated into changes in the measured spectrum. This reduces the problem to modeling how the spectrum changes with TX position, which can be viewed as a 2D interpolation task over space. In the RX-side moving case, the WRF is static, and the RX samples the spectrum from different positions. This corresponds to a 3D reconstruction problem, for which 3D Gaussian modeling is a natural fit. However, since the WRF itself does not change, the variation in the observed spectrum is entirely due to the change in the field-to-spectrum mapping, which is a function of the RX position. This perspective allows the task to be reformulated as a 2D interpolation problem over RX positions, consistent with the TX-side moving case.
3.3.2 Problem Definition
We now formally define the task. To begin, we describe the two settings considered under single-sided transceiver mobility:
TX-side moving scenario. The TX moves across a set of known positions
$\mathcal{P}=\{\boldsymbol{p}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N_{s}},$
RX-side moving scenario. The TX emits WRF signals from a fixed position $\boldsymbol{p}\in\mathbb{R}^{3}$, while the RX moves across a set of known positions
$\mathcal{Q}=\{\boldsymbol{q}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N_{s}},$
$\{(\boldsymbol{s}_{i},\mathbf{A}_{i})\}_{i=1}^{N_{s}},\quad\boldsymbol{s}_{i}\in\begin{cases}\mathcal{P},&\text{TX-side moving},\\ \mathcal{Q},&\text{RX-side moving},\end{cases}$
Figure 4: 2DGS formulation and spectrum rasterization.